# More than ‘Left and Right’: Revealing Multilevel Online Political Selective Exposure

### v1

@@ 4.2 Multi-scale Community Detection
We then apply a multi-scale community detection approach to the projected influencer graph $\mathcal{G}^{\mathcal{S}}$. Community detection is the process of identifying groups of nodes that are more densely connected to each other than to the rest of the network. It is common to apply community detection methods in studying phenomena like selective exposure, echo chambers, and polarization online [51, 26, 5, 52]. However, most studies detect communities using single-scale approaches, such as modularity optimization [53]. But real-world social networks are usually multi-scale and hierarchical [42, 43]. Multi-scale community detection approaches allow for the discovery of communities at different resolution scales. This flexibility enables selecting the scales closest to the real situation among all the experimental scales [54].
@@
A variety of techniques have been designed for multi-scale community detection, such as methods with adjustable resolution parameters (e.g., [55, 56, 44]) and methods which automatically detect relevant scales (e.g., [57, 58, 45]). Here, we use Markov Stability[44, 59] with automatic scale selection[45] and optimization using the Leiden algorithm [60]. This method finds communities as groups of nodes that best retain the flow of a random walk process. For each time scale of the process, communities of different scales are found, with longer times corresponding to larger scales. Optimal scales are found as minima of the variation of information of an ensemble partitions computed at each scale [59]. The process results in seven optimal scales. We discard the two first scales with 1767 and 1170 communities, respectively, as they have very low granularity, and more than 80 percent of the communities only include one node. Finally, we have five levels for analysis, with 46, 14, 8, 3, and 2 communities respectively. More information about running community detection can be found in section 4 in SI.
For each scale $s$, the community detection yields a partition of the influencer network $\mathcal{G}^{\mathcal{S}}$ in $K_{s}$ non-overlapping communities $\{C^{s}_{i}\}_{i=1}^{K_{s}}$. By projecting the partitions on the consumer network $\mathcal{G}^{\mathcal{C}}$, we obtain, possibly overlapping, communities of consumers where two nodes are in the same community if they follow influencers belonging to the same community. This overlapping partitioning of the consumer network can also be seen as a hypergaph, where for each community $C_{i}^{s}$ at scale $s$ in the influencer network, corresponds a hyperedge, $E_{i}^{s}$ in the consumer network defined as the set of users following at least one influencer in $C_{i}^{s}$. At scale $s$, $\{E^{s}_{i}\}_{i=1}^{K_{s}}$ is therefore the set of $K_{s}$ hyperedges of the consumer network mapping the community partition $\{C^{s}_{i}\}_{i=1}^{K_{s}}$ of the influencer network.
4.3 Measurements of Selective Exposure Indices
We introduce five indices – Community Overlap, Identity Diversity, Information Diversity, Structural Isolation, and Connectivity Inequality to measure individuals’ selective exposure patterns at different levels.
@@ 4.3 Measurements of Selective Exposure Indices
Community Overlap. Community Overlap is measured for each consumer node and at each level and corresponds to the number of influencer communities a consumer belongs to. As shown in Fig.4(a), each consumer node can follow influencers either in a single community (annotated in green) or from multiple communities (annotated in purple).
@@
At scale $s$, for a consumer node $v\in\mathcal{C}$, let $I(v,E_{i}^{s})$ be an indicator function that is 1 if node $v$ belongs to the hyperedge $E_{i}^{s}$, and 0 otherwise. The number of hyperedges nodes $v$ belongs to, i.e., the number of influencer communities that the users it follows belong to, is expressed as
@@
Identity Diversity. Identity Diversity measures the diversity of annotated identities of members within a influencer community. We annotate identities for influencers on four dimensions: Political Ideology, Campaign Support, Social Identity, and Account Type. We only measure the diversity of labels on the Political Ideology dimension as it is the most direct indicator of political identity, and we deal with the unlabeled accounts with probability assignment, which is introduced below.
@@
Where $L$ is the number of ideological labels (here:“Left”, “Right” and “Center”), $n_{\ell}$ is the number of influencers of label $\ell$ in community $C_{i}^{s}$, and $N$ is the total number of influencers in community $C_{i}^{s}$. As not all the influencers are labeled, we apply two approaches to deal with the missing values: in one approach, we assign probabilities of labels to the unlabeled influencer accounts based on the proportions of labels present in each local community; in the other approach, we only consider the influencer accounts that are labeled. The two approaches yield similar results (see section 5 in SI).
Information Diversity. We also use the Gini-Simpson Diversity Index to measure the diversity of website domains posted by the influencers in that community, which is defined as Information Diversity. The Information Diversity, $\text{InfD}(C_{i}^{s})$ of community $C_{i}^{s}$ at scale $s$ is used to measure exposure at the information level and is computed with (2), with $L$ as the number of different website domains, $n_{\ell}$ the number of times a website from domain $\ell$ is shared by an influencer of community $C_{i}^{s}$, and $N$ is the total number of links shared by influencers in community $C_{i}^{s}$.
This index measures how diversified the information sources are witnessed by individual consumers who follow that community instead of just the diversity of influencers’ profiles. To achieve better reliability, we remove the communities that share fewer than 100 website domains and have less than 50 percent of influencer sharing links (see section 5 in SI).
Structural Isolation. The Normalized Cut measures the strength of the connection between two disjoint sets of nodes splitting a network in two and is a popular measure in image segmentation and data clustering [62]. We employ it to measure the degree of isolation of a community by computing the normalized cut between the nodes inside and outside the community, defined as Structural Isolation. For a community $C_{i}^{s}$, at scale $s$, the normalized cut takes into account the internal and external connectivity of $C_{i}^{s}$ and is computed as
@@
where $c_{i}^{s}$ is the size of the cut, i.e., the sum of the weights of the edges crossing the boundary of the community $c_{i}^{s}=\sum_{u\in C_{i}^{s},v\notin C_{i}^{s}}w_{uv}$, $m_{i}^{s}$ is the sum of the weights of the edges inside of the community, i.e. $m_{i}^{s}=\sum_{u,v\in C_{i}^{s}}w_{uv}$, and $m$ is the sum total of the edge weights in the influencer network, $m=\sum_{u,v\in\mathcal{S}}w_{uv}$.
Connectivity Inequality. We measure the degree of inequality of the degree weights of influencers, which represents the number of shared consumers with other influencers, in a community with the Gini Index[63]. It measures the extent to which the political information consumers are mainly exposed to certain influential political accounts or distribute equal attention to a broad set of influencers and is computed as
@@
where $N_{i}^{s}$ is the number of nodes in the community $C_{i}^{s}$, $k_{u}=\sum_{v\in C_{i}^{s}}w_{uv}$ is the weighted internal degree of node $u$ and $k_{i}^{s}\right>$ is the average weighted internal degree of the nodes in community $C_{i}^{s}$.
4.4 Variable Dimension Reduction and Regression Analysis
@@
We perform a regression analysis to explore factors related to the five indices of selective exposure patterns. The independent variables are consumers’ individual attributes. To examine a wide range of consumers’ attributes, we construct 10 groups of 189 variables obtained from the National Survey in Brazil, including Demographics, News Consumption, Political Communication, Political Identification, Political Engagement, Perceptions of Incivility, Perceptions of Disinformation, Authority Trust, Populism, and Attitudes to Democracy. These groups encompass individuals’ social, political, and media attributes. We then reduce the dimensions of variables in each group by applying PCA, retaining only one variable for each significant dimension and 25 variables in total (see section 6 in SI).
@@
The dependent variables are measurements of five indices of selective exposure patterns. Except for Community Overlap, the other four indices — Identity Diversity, Information Diversity, Structural Isolation, and Connectivity Inequality — are initially measured based on influencer communities. We transform these four indices onto individual levels by calculating the average values of influencer communities each consumer interacts with, weighted by the number of links between consumers and influencers in each community.

### v2

@@ Introduction
We then apply a multi-scale community detection approach to the projected influencer graph $\mathcal{G}^{\mathcal{S}}$. Community detection is the process of identifying groups of nodes that are more densely connected to each other than to the rest of the network. It is a common strategy to study phenomena like selective exposure, polarization, and echo chambers [59, 27, 5, 60]. However, most studies detect communities using single-scale approach, such as modularity optimization [45]. But real-world social networks are usually multi-scale and hierarchical, for example, smaller class groups are embedded within the broader structure of a school [36, 37]. Multi-scale community detection approaches allow for the discovery of communities at different resolution scales. This flexibility also enables selecting the scales closest to the real situation among all the experimental scales, depending on the research goals [38].
@@
A variety of techniques have been designed for multi-scale community detection, such as methods with adjustable resolution parameters (e.g., [39, 40, 35]) and methods which automatically detect relevant scales (e.g., [41, 42, 8]). Here, we use Markov Stability[35, 7] with automatic scale selection[8] and optimization using the Leiden algorithm [46]. We note that our approach is applicable to the results of any multiscale community detection approach. The rationale of the Markov Stability method is to identify communities as groups of nodes where a random walk remains confined over varying time scales, thus revealing community structures at different resolutions. We give a detailed mathematical explanation of the Markov Stability method in section 5 in Supplementary Information.
Multidimensional annotation of identities
We then annotate four identity attributes of political influencer nodes in the network: political ideology, campaign support, social identity, and account type. The data are manually annotated based on the self-presentation of political and social identities in users’ Twitter/X profiles, combined with a formal list of politicians and parties participating in the election. It is shown that users who disclose their political identities publicly are generally more active than those who do not [63]. Almost all political influencers have profile descriptions.
Political ideology is the most traditional and straightforward indicator of individuals’ political opinions and comprises categories {Left, Right, Center}. During the 2022 Brazilian Presidential Election, Lula and Bolsonaro are the two candidates who represent leftist and rightist leaders, respectively. Nevertheless, it is not necessarily the case that all left-wing individuals support Lula and all right-wing individuals support Bolsonaro, due to increasing cleavages on issues and socio-cultural factors, even though there is still significant overlap. To differentiate between these situations, we also annotate them as ”Pro-Lula” or ”Pro-Bolsonaro” if they support certain candidates during the election campaign.
Many users also indicate their social identities, as cultural issues have become increasingly related to politics. We identify four social identities most frequently revealed in the profiles: Women (in the context of feminism or supporting women’ rights), Religious, Black, and LGBTQ. For religious, Black, and LGBTQ identities, annotation is based on either users’ self-identification with the group or their expression of support for its rights. Note that, unlike the fixed options of political ideology, social identity encompasses more categories; accordingly, we include only those observed in our samples. The same situation applies to account type, which indicates what the accounts represent. We consider media (including media outlets and journalists), politicians, and individual opinion leaders, as they are the most common accounts that produce political information.
As not all political influencers reveal their identities at all dimensions, we assign those unrevealed profiles as “Unlabeled”. The annotation results on political ideology, campaign support, social identity, and account types for influencers in communities of all scale levels are included in section 4 in Supplementary Information.
Measurements of selective exposure indices
We introduce five indices – Community Overlap, Identity Diversity, Information Diversity, Structural Integration, and Connectivity Inequality to measure individuals’ selective exposure patterns at different levels. We consider a followship bipartite network between political influencers and survey respondents, and project the network onto both the survey participants’ side and the political influencers’ side. From the projection on the survey participants, we capture the number of influencer communities in which they participate. From the projection on the political influencers, we detect communities and identify the characteristics of influencer communities followed by each survey participant. In this context, we introduce the following measures:
Community overlap.
This index is derived from the network’s projection on the survey participants and reflects the number of influencer communities they engage with. It captures how online users are divided into fragmented communities in which individuals primarily interact with others who share similar interests, and how they could act as bridges between multiple communities. This index is based on the phenomenon of cyberbalkanization introduced by Alstyne and Brynjolfsson in their 2005 study [7] which showed that some internet users are more inclined to have diverse interests and engage with multiple communities, while others prefer to remain within a single community. Survey users who connect with multiple communities have a higher potential to act as bridgers—individuals who consume information from a variety of sources [49]. To quantify survey respondents’ capacity to bridge multiple communities, we introduce the Community Overlap Index.
@@
Community Overlap is measured for each consumer node and at each level and corresponds to the number of influencer communities a consumer belongs to. Each consumer node can follow influencers either in a single community or from multiple communities.
@@
At scale $s$, for a consumer node $v\in\mathcal{C}$, let $I(v,E_{i}^{s})$ be an indicator function that is 1 if node $v$ belongs to the hyperedge $E_{i}^{s}$, and 0 otherwise. The number of hyperedges node $v$ belongs to, i.e., the number of influencer communities that the users it follows belong to, is expressed as
@@
Identity diversity.
This index is measured from the network’s projection on the political influencers and indicates the diversity of identities among the influencers engaged by survey participants. According to social identity theory and social categorization theory, people may act based on inter-group behaviors arising from a sense of belonging to certain social groups [50, 51]. It can be applied to the political domain, where political identities advance social identities, such as national identities, partisan affiliations, and political ideologies. [52]. The diversity of influencers’ political identity in a community is measured by the Identity Diversity Index.
We annotate identities for influencers on four dimensions: Political Ideology, Campaign Support, Social Identity, and Account Type, as indicated in their profile descriptions. Influencers who strongly exhibit certain attributes may lead individuals to filter into specific community structures. However, not all influencers disclose identities in all dimensions. To address unlabeled accounts, we apply probability assignment, as described below (also see the percentages of unlabeled identities in Figs E–H in Supplementary Information). We only measure the diversity of labels on the Political Ideology dimension as it is the most direct indicator of political identity.
@@
Where $L$ is the number of ideological labels (here:“Left”, “Right” and “Center”), $n_{\ell}$ is the number of influencers of label $\ell$ in community $C_{i}^{s}$, and $N$ is the total number of influencers in community $C_{i}^{s}$. As not all the influencers are labeled, we apply two approaches to deal with the missing values: in one approach, we assign probabilities of labels to the unlabeled influencer accounts based on the proportions of labels present in each local community; in the other approach, we only consider the influencer accounts that are labeled. The two approaches yield similar results (see section 6 in Supplementary Information).
Information diversity.
Likewise, this index is computed based on the communities found on the network’s projection on the influencers. Instead of being exposed to certain political identity groups, ordinary users may focus on the information retrieved from specific communities. People may have a preference for certain information sources or opinions due to personal interests or confirmation bias [53, 54]. Here, we measure the extent to which individuals are exposed to diverse information using the website domain links shared by political influencers.
Similar to Identity Diversity, we use the Gini-Simpson Diversity Index to measure the diversity of website domains shared by the influencers in that community, which is defined as Information Diversity. The Information Diversity, $\text{InfD}(C_{i}^{s})$ of community $C_{i}^{s}$ at scale $s$ is used to measure exposure at the information level and is computed with (2), with $L$ as the number of different website domains, $n_{\ell}$ the number of times a website from domain $\ell$ is shared by an influencer of community $C_{i}^{s}$, and $N$ is the total number of links shared by influencers in community $C_{i}^{s}$.
This index measures how diversified the information sources are witnessed by individual consumers who follow that community instead of just the diversity of influencers’ profiles. To achieve better reliability, we remove the communities that share fewer than 100 website domains and have less than 50 percent of influencer sharing links (see section 6 in Supplementary Information).
Structural integration.
This is a structural measure of the influencer communities engaged by survey participants. Each influencer community occupies a distinctive structural position within the network. Some communities are relatively isolated from others, while some are more integrated into the overall structure. In the co-following influencer network, a more isolated community indicates that it caters to a more targeted consumer group, whereas greater integration suggests exposure to a broader audience.
To quantify this, we introduce the Structural Integration Index. This index is derived from the Normalized Cut, a widely used metric in image segmentation and graph partitioning that evaluates the connectivity between two disjoint subsets of nodes in a network [56]. We employ the Normalized Cut to assess the degree of integration of a given community by comparing its internal and external connectivity. For a community $C_{i}^{s}$ at scale $s$, the Structural Integration is defined as:
@@
where $c_{i}^{s}$ is the size of the cut, i.e., the sum of the weights of the edges crossing the boundary of the community $c_{i}^{s}=\sum_{u\in C_{i}^{s},v\notin C_{i}^{s}}w_{uv}$, $m_{i}^{s}$ is the sum of the weights of the edges inside of the community, i.e. $m_{i}^{s}=\sum_{u,v\in C_{i}^{s}}w_{uv}$, and $m$ is the sum total of the edge weights in the influencer network, $m=\sum_{u,v\in\mathcal{S}}w_{uv}$. The Structural Integration Index reaches a minimum value of 0 and a maximum value of 2 for two degenerate cuts: the empty cut, and the cut containing all edges. In this work, all the values of the index were between 0 and 1. Lower values indicate stronger isolation with the rest of the network, and larger values reflect greater integration, meaning the community has relatively more or stronger connections to other parts of the network. This measure enables us to evaluate whether survey participants are primarily exposed to influencers from structurally peripheral communities or from those that are more centrally embedded in the network.
Connectivity inequality.
This index is also based on the influencer communities but focuses on their internal structures. Some communities exhibit an unequal distribution of (weighted) degrees, with certain nodes having a larger number of connections and potentially playing a more central role, while others display a more even distribution.
We quantify the inequality in the distribution of influencer degree weights—defined as the number of shared survey consumers each influencer has with others within a community—using the Gini Index [57]. It captures the skewness of the degree distribution. A highly skewed distribution may arise due to the presence of “super-influencers”, such as prominent politicians, who share survey consumers with many other influencers. The metric is computed as
@@
where $N_{i}^{s}$ is the number of nodes in the community $C_{i}^{s}$, $k_{u}=\sum_{v\in C_{i}^{s}}w_{uv}$ is the weighted internal degree of node $u$ and $\left<k_{i}^{s}\right$> is the average weighted internal degree of the nodes in community $C_{i}^{s}$.
Finally, all influencer community-level indices—Identity Diversity, Information Diversity, Structural Integration, and Connectivity Inequality—are aggregated at the user level for the regression analysis by calculating the weighted average across all influencer communities with which the individual is connected.
Specifically, let $\mathcal{C}_{u}$ denote the set of influencer communities followed by survey participant $u$, and let $w_{uC_{j}}$ represent the number of influencers in community $C_{j}\in\mathcal{C}_{u}$ followed by user $u$. For each index $I\in\{\text{IdD},\text{InfD},\text{SI},\text{CI}\}$, the aggregated individual-level score for user $u$ is computed as a weighted average of the community-level values:
$I_{u}=\frac{\sum_{C_{j}\in\mathcal{C}_{u}}w_{uC_{j}}\cdot I(C_{j})}{\sum_{C_{j}\in\mathcal{C}_{u}}w_{uC_{j}}}$ (5)
This step ensures that communities with more connections to a given user contribute proportionally more to their overall exposure, reflecting the strength of affiliation and potential information influence. In contrast, the Community Overlap Index ($\text{CO}^{s}(v)$) is inherently computed at the individual level based on the number of influencer communities followed, and thus does not require further aggregation.
Variable dimension reduction and regression analysis
@@
Having measured the patterns of consumers’ selective exposure, we continue to examine the association of ten groups of individual attributes obtained from the survey: Demographics, News Consumption, Political Communication, Political Identification, Political Engagement, Perceptions of Incivility, Perceptions of Disinformation, Authority Trust, Populism, and Attitudes toward Democracy, with the five selective exposure indices (see section 7 in Supplementary Information for detailed description of the ten groups of individual attributes).
We perform a regression analysis to explore factors related to the five indices of selective exposure patterns. The independent variables are consumers’ individual attributes. To examine a wide range of consumers’ attributes, we construct 10 groups of 189 variables obtained from the National Survey in Brazil, including Demographics, News Consumption, Political Communication, Political Identification, Political Engagement, Perceptions of Incivility, Perceptions of Disinformation, Authority Trust, Populism, and Attitudes to Democracy. These groups encompass individuals’ social, political, and media attributes. We then reduce the dimensions of variables in each group by applying PCA, retaining only one variable for each significant dimension and 25 variables in total (see section 7 in Supplementary Information).
@@
The dependent variables are measurements of five indices of selective exposure patterns. As mentioned before, except for Community Overlap, the other four indices—Identity Diversity, Information Diversity, Structural Integration, and Connectivity Inequality—are initially measured based on influencer communities. We transform these four indices onto individual levels by calculating the average values of influencer communities each consumer interacts with, weighted by the number of links between consumers and influencers in each community.